# Evolutionary Prediction Games

### v1

@@ 3 Evolutionary Prediction Games
We note that a Nash equilibrium exists for any population game (Sandholm, 2010, Thm. 2.1.1). In relation to fairness, we make the following connection (see Section B.2.2):
Proposition 3.2.
Let $F({\bm{p}})$ be an evolutionary prediction game induced by a learning algorithm $\mathcal{A}$, and let ${\bm{p}}^{*}$ be a Nash equilibrium. Then $\mathcal{A}$ satisfies the overall accuracy equality criterion in expectation when trained on $D_{{\bm{p}}^{*}}$.
4 Survival of the Fittest
4.1 Warm-up: Training Once
We begin with a characterization of evolutionary outcomes for a simple setting in which the classifier is trained only once on the initial distribution, and henceforth kept fixed. This is useful as a first step since, given the learned $h$, outcomes can be attributed solely to changes in the population.
Proposition 4.1.
@@ 4.1 Warm-up: Training Once
    $\mathrm{acc}_{{\bm{p}}^{*}}(h^{0})\geq\mathrm{acc}_{{\bm{p}}^{0}}(h^{0})$
@@
    ${\bm{p}}^{*}$ is supported on $k^{*}=\operatorname*{arg\,max}_{k}\mathrm{acc}_{k}(h^{0})$
@@
    ${\bm{p}}^{*}$ is stable
Proof in Section B.3. Proposition 4.1 makes two important statements. First, from the perspective of the prediction system operator, overall accuracy only improves over time as a result of population dynamics. Since only group proportions ${\bm{p}}$ change, this improvement is due to an increase in $p_{k}$ for higher-accuracy groups, and a decrease for lower-accuracy groups. Thus, average accuracy increases—though not because of an actual ‘improvement’ in the model, but rather, due to the distribution becoming easier to predict.
Second, in the limit, this effect is driven to extreme, and a single group $k^{*}$—the one having the highest initial accuracy under $h^{0}$—will dominate at equilibrium ${\bm{p}}^{*}$. In evolution, this is known as the competitive exclusion principle (Hardin, 1960), which states that if multiple species compete over the same resources, then even a minimal advantage of one species over the others results in exclusive survival of the fittest species, and extinction of the rest.Our result suggests that this strong form of ‘survival of the fittest’ also applies to group dynamics driven by a fixed classifier.
@@
Although our results forecast that only the fittest will survive, it is important to remember that this describes outcomes in the limit. A more practical takeaway is that, as time progresses, the general tendency of the system will be for $p_{k^{*}}$ to gradually grow on account of others. If we wish to prevent eventual domination, then we should seek ways to intervene along the way. One possible approach is to retrain $h$, in hopes that this will lead to different outcomes: note that while accuracy under $h^{0}$ improves over time, the final $\mathrm{acc}_{{\bm{p}}^{*}}(h^{0})$ is by no means guaranteed to be the maximal attainable accuracy. Thus, working to continually improve accuracy should affect the eventual population. We now turn to study population dynamics under retraining.
4.2 Retraining Optimal Classifiers
Figure 3: Evolutionary dynamics for two groups, induced by an optimal linear classifier (Section 4.2). (Left) Data distributions in feature space. Dashed line represents the optimal linear classifier $h_{\bm{p}}$ for the uniform mixture ${\bm{p}}=(0.5,0.5)$. (Center) Evolutionary prediction game induced by an optimal linear classifier. The game has two stable equilibria with single-group dominance, and one unstable coexistence equilibrium. Overall population accuracy $\mathrm{acc}_{\bm{p}}(h_{\bm{p}})$ is convex, and maximized at the stable equilibria. (Right) Replicator dynamics induced by the population game, for various initial conditions. Populations evolve towards states with single-group dominance.
To understand how populations and classifiers evolve jointly, we start by characterizing evolutionary outcomes for an ‘ideal’ setting in which the deployed classifier is optimal:
Definition 4.2 (Optimal classifier).
Let $\mathcal{H}$ be a hypothesis class, and let $D$ be a data distribution. An optimal classifier w.r.t. $\mathcal{H}$ is a minimizer of the 0-1 loss in expectation over $D$:
Working with optimal classifiers abstracts away the complexities due to estimation and approximation errors of practical learning algorithms (i.e., algorithms that optimize proxy loss functions over finite data). Nonetheless, analyzing a dynamical system that is driven by the argmin of an expected objective remains quite challenging. We focus on settings in which $h^{\mathrm{opt}}$ exists, and denote by $\mathcal{A}^{\mathrm{opt}}({\bm{p}})$ the (theoretical) learning algorithm which returns the optimal classifier with respect to $D_{\bm{p}}$.222 Bayes-optimal classifiers $h^{\mathrm{Bayes}}(x)=\operatorname*{arg\,max}_{y}p(y\,|\,x)$ are a special case of this definition, for the hypothesis class $\mathcal{H}=\mathcal{Y}^{\mathcal{X}}$. In this case, the fitness function becomes $F_{k}({\bm{p}})=\mathrm{acc}_{k}\left(\mathcal{A}^{\mathrm{opt}}({\bm{p}})\right)$. We assume that fitness functions are differentiable, and that that ties are broken consistently if multiple optimal classifiers exist.
Our central theoretical result characterizes the possible evolutionary outcomes induced by optimal classifier retraining:
Theorem 4.3.
Let $\mathcal{H}$ be a hypothesis class, and denote by $\mathcal{A}^{\mathrm{opt}}$ the optimal learning algorithm w.r.t. $\mathcal{H}$. Fix an initial state ${\bm{p}}^{0}$, and assume that at each population state ${\bm{p}}$ the optimal classifier $h_{\bm{p}}\sim\mathcal{A}^{\mathrm{opt}}({\bm{p}})$ is deployed. Then it holds that:
  1. 1.
    Stable equilibria always exist. For all stable equilibria:
    • ${\left|{\operatorname*{support}({\bm{p}}^{*})}\right|}=1$
    • $\mathrm{acc}_{{\bm{p}}^{*}}\left(h_{{\bm{p}}^{*}}\right)\geq\mathrm{acc}_{{\bm{p}}^{0}}\left(h_{{\bm{p}}^{0}}\right)$
  2. 2.
    Unstable equilibria with ${\left|{\operatorname*{support}({\bm{p}}^{*})}\right|}>1$ may exist.
Proof in Section B.4. To prove the claim, we leverage basic structure which is revealed despite the complexity of the learning problem. Our main technical tool is the framework of potential games, which are population games that can be expressed as the gradient of a potential function. We first show that the expected accuracy of any fixed classifier is linear in ${\bm{p}}$. Then the core of the proof is a convexity argument: The optimality of $\mathcal{A}$ implies that $\mathrm{acc}_{\bm{p}}(h_{\bm{p}})$ is convex as a pointwise maximum of linear functions, with gradients given by marginal accuracies. From this we identify $\mathrm{acc}_{\bm{p}}(h_{\bm{p}})$ as the potential function for the game, and leverage the known correspondence between Nash equilibria of population games and extremal points of potential functions. Finally, the stability of single-population states follows from the fact that maximal points of convex functions over the simplex are attained at the vertices.
@@ 4.2 Retraining Optimal Classifiers
Theorem 4.3 suggests multiple equilibria can exist, and distinguishes between two types. If dynamics converge to a stable equilibrium ${\bm{p}}^{*}$, then accuracy improves, but this is again through competitive exclusion, since $p^{*}_{k}>0$ for only one group. However, in contrast to training once, the surviving group is no longer predetermined, and can depend on context (e.g., initial state) and choices (e.g., model class). The more critical difference, however, is that retraining enables coexistence, in the form of possible equilibria in which multiple groups survive. That is, there may exist population states ${\bm{p}}^{*}$ such that $p^{*}_{k}>0$ for several groups $k$, and for which this remains to hold when the classifier is retrained. The crux, unfortunately, is that such points are inherently unstable, in the sense that any small perturbation of ${\bm{p}}^{*}$ will push dynamics away. Hence, under plausible dynamics, and for retraining with optimal classifiers, it is most unlikely that dynamics will converge to such states.
5 Avenues for Coexistence
Figure 4: Graphical illustration of coexistence results. Top row presents data distributions, bottom row shows the corresponding prediction games. (Left) Stable and beneficial coexistence induced by the use of a proxy loss (Theorem 5.1). (Center) Stable coexistence induced by the use of memorization (Theorem 5.2). (Right) Beneficial coexistence induced by learning from a finite dataset (Theorem 5.3).
Intuitively, optimal classifiers drive dynamics away from mixture equilibria because the learning objective is fully aligned with, and so reinforces, fitness. But in practice we must resort to optimizing proxy empirical objectives, which can break alignment. This raises two important questions: when will mixture equilibria be stable, and can we make them such? For the first question, we show that each of the aspects which enable learning in practice—namely the use of proxy losses, finite data access, or memorization—can create conditions for coexistence to organically materialize. This entails coexistence that is beneficial (w.r.t. accuracy) stable, or both. For the second, we present preliminary steps.
5.1 Proxy Loss
A common approach to overcoming the difficulty due to the discontinuity of the 0-1 loss is to replace it with a convex surrogate defined over predictive scores. This enables optimization, but introduces bias: for example, instead of a uniform penalty of 1, misclassified points are penalized relative to their distance from the decision boundary of $h$. Our next result shows that this limitation enables coexistence if the two groups have opposing biases, which balance at equilibrium. The construction relies on the Soft-SVM algorithm which uses the hinge loss with $\ell_{2}$ regularization.
Theorem 5.1.
For the Soft-SVM algorithm, there exist data distributions that induce an evolutionary prediction game with a mixture equilibrium that is both stable and beneficial.
Proof in Section B.5, and the construction is outlined in Figure 4 (Left). The main idea of the construction is to leverage the bias of Soft-SVM against the minority class in imbalanced settings. For each group alone, minimizing the proxy loss biases the learned classifier against the minority class, which is suboptimal. However, when both groups coexist, these biases cancel out. The surprising property is that when one group grows slightly (e.g., $p_{1}=0.5+\epsilon$), it loses more from the fact that the other group shrinks (e.g., $p_{2}=0.5-\epsilon$) than it gains from its own growth. This symbiotic relation leads to stable and beneficial coexistence. In Section B.5.3, we expand on this idea, showing that varying the SVM regularization parameter leads to bifurcations in which the game transitions between 3, 1, and 5 equilibria.
5.2 Memorization
Another practical approach to loss minimization is to use a class of models that fit the training data perfectly and then interpolate. Notable examples $k$-NN and overparameterized neural networks (e.g., Zhang et al., 2021). Memorization leads to zero loss on the training set by definition, but often at the price of sensitivity to noise; as a concrete example, consider how 1-NN predictions change with the addition of a single point. The problem is that even in the limit, 1-NN cannot express the true $p(y|x)$, and remains sensitive to label noise. Our next result shows this can lead to coexistence:
Theorem 5.2.
Proof in Section B.6, and the construction is illustrated in Figure 4 (Center). The key property leveraged by this proof is asymmetric label noise. There are two groups defined symmetrically. Each group is composed of a majority and minority classes, and the majority class has label noise. For each group alone, the 1-NN algorithm makes perfect predictions on its minority class, and imperfect predictions on the majority class. Since groups are defined symmetrically, training on one group alone leads to better accuracy for the other—leading to stable coexistence. Section B.6.2 shows that under $k>1$, stable coexistence can also be beneficial.
5.3 Finite data
All practical learning algorithms must rely on a finite sample of data points for training a classifier. Learnability (in the PAC sense) implies that as data size grows, the sample becomes increasingly representative of the distribution. But the rate at which variance diminishes need not be equally across all regions of a distribution. In a sense, this means that one region can be learned ‘faster’ (i.e., with less data) than another. Our final example exploits this to show how coexistence can arise from statistical considerations.
Theorem 5.3.
Proof in Section B.7, and the construction is illustrated in Figure 4 (Right). The key property we leverage here is asymmetric class variance. In each group, one class has high variance, and the other class has low variance. When the dataset is finite (of size $n$), this creates bias against the high-variance class, and thus convergence towards the optimal classifier at a linear rate ($O(n^{-1})$). When the two groups coexist, this bias is balanced, and the rate of convergence is exponential—leading to beneficial coexistence. The crux in the construction is that while coexistence is beneficial, it is unstable. Generally we suspect that stability is difficult to obtain through pure statistical means. But this does not mean that it is generally unattainable, as we show next.

### v2

@@ 3 Evolutionary Prediction Games
This also applies to evolutionary prediction games; e.g., the game in Figure 2 (Center) has three Nash equilibria at ${\left\{{(1,0),(0.5,0.5),(0,1)}\right\}}$, and all population games have at least one equilibrium [43].
For evolutionary prediction games, Nash equilibria have a tight connection to fairness (Appx. D.3.2):
Proposition 1.
Let $F({\bm{p}})$ be an evolutionary prediction game induced by a learning algorithm $\mathcal{A}$, and let ${\bm{p}}^{*}$ be a Nash equilibrium. Then $\mathcal{A}$ satisfies the overall accuracy equality fairness criterion in expectation when trained on $D_{{\bm{p}}^{*}}$.
Induced population dynamics.
Formally, our theoretical analysis encodes evolutionary dynamics as $\dot{\bm{p}}=V_{F}({\bm{p}})$, where $\dot{\bm{p}}$ is the time derivative of ${\bm{p}}$, and $V_{F}:\Delta^{K}\to T\Delta^{K}$ is a tangent vector field induced by the game $F({\bm{p}})$. Our results require only mild structural assumptions on $V_{F}$, namely: (i) $V_{F}$ is continuous; (ii) the direction of flow is aligned with fitness (i.e., $V_{F}({\bm{p}})\cdot F({\bm{p}})>0$ whenever $V_{F}\neq 0$, a property known as positive correlation [43]); (iii) fixed points of the dynamical system coincide with the Nash equilibria of $F({\bm{p}})$. All are satisfied by the user-level dynamics highlighted in Section 1 and Appendix C, where they emerge as group-level properties in the large-population limit.111 We note that imitative dynamics may introduce additional fixed points corresponding to restricted equilibria of the game, but these do not affect our main theoretical results due to their instability. See Section D.2. Moreover, many of them converge towards the canonical replicator equation, $\dot{p}_{k}=p_{k}\left(F_{k}({\bm{p}})-\bar{F}({\bm{p}})\right)$, where $\bar{F}({\bm{p}})=\sum_{k^{\prime}}p_{k^{\prime}}F_{k^{\prime}}$ is the average fitness across the population [48]. We note that time-scales of convergence are determined (and adjustable) by the choice of $V_{F}$.
Heterogeneous fitness.
Finally, in some settings, incentives or outside alternatives may vary across groups. This may be captured using the notion of retention functions [19], which translate marginal prediction accuracy to the tendency to remain engaged. In such cases, the transformed fitness function becomes $F_{k}({\bm{p}})=\nu_{k}\left(\mathbb{E}_{h_{\bm{p}}}{\left[{\mathrm{acc}_{k}(h)}\right]}\right)$, where $\nu_{k}:[0,1]\to\mathbb{R}$ is a group-dependent, continuous strictly-increasing function. A simple example is $\nu_{k}(x)=x-b_{k}$ (for $b_{k}>0$) which allows groups to differ in the minimum accuracy their users require to keep using the system. All of our results apply to any set of strictly monotone $\nu_{k}$ for $K=2$, and to positive affine $\nu_{k}$ with per-group offsets $b_{k}$ for $K>2$. For clarity we state our main results for $\nu_{k}(x)=x$, but also discuss them for broader $\nu$.
4 Competitive Exclusion Under Oracle Classifiers
To understand how populations and classifiers evolve jointly, we start by characterizing natural selection in an ‘ideal’ setting, in which a classifier is trained repeatedly using unlimited resources:
Definition 1 (Oracle classifier).
Let $\mathcal{H}$ be a hypothesis class, and let $D$ be a data distribution. An oracle classifier with respect to $\mathcal{H}$ is a minimizer of the 0-1 loss in expectation over $D$:
Oracle classifiers represent ERM classifiers in the population limit. This regime abstracts away the complexities due to estimation and approximation errors of practical learning algorithms (i.e., algorithms that learn from finite data in reasonable time). Nonetheless, analysis of the dynamics remains challenging due to the $\mathrm{argmin}$ operator. We focus on settings in which $h^{\mathrm{opt}}$ exists, and denote by $\mathcal{A}^{\mathrm{opt}}(D)$ the (theoretical) learning algorithm which returns the oracle classifier with respect to $D$222 The Bayes-optimal classifier $h^{\mathrm{Bayes}}(x)=\operatorname*{arg\,max}_{y}p(y\,|\,x)$ is a special case of this definition for $\mathcal{H}=\mathcal{Y}^{\mathcal{X}}$. . We assume that fitness functions are continuous, and tie-breaking is consistent. To simplify presentation, we first assume that each group has a distinct oracle classifier, and then address the general case. Our central result characterizes natural selection induced by oracle classifier retraining:
Theorem 1.
Let $\mathcal{H}$ be a hypothesis class, and denote by $\mathcal{A}^{\mathrm{opt}}$ the oracle learning algorithm with respect to $\mathcal{H}$. Assume that at each population state ${\bm{p}}$ the oracle classifier $h_{\bm{p}}\sim\mathcal{A}^{\mathrm{opt}}(D_{\bm{p}})$ is deployed. Then it holds that:
@@ 4 Competitive Exclusion Under Oracle Classifiers
    Accuracy: Overall accuracy increases over time, $\tfrac{\mathrm{d}}{\mathrm{d}t}\mathrm{acc}_{{\bm{p}}}(h_{{\bm{p}}})\geq 0$.
@@
    Stability: A stable equilibrium always exists, and there can be multiple such equilibria.
@@
    Competitive exclusion: For all stable equilibria, ${\left|{\operatorname*{support}({\bm{p}}^{*})}\right|}=1$.
  4. 4.
    Coexistence: Equilibria with ${\left|{\operatorname*{support}({\bm{p}}^{*})}\right|}\geq 2$ may exist, but are unstable.
Proof sketch.
The proof leverages structural properties that emerge despite the complexity of the learning problem. Our main technical tool is the framework of potential games, which are population games that can be expressed as a gradient of a scalar potential function defined over the simplex. We start by showing that the expected accuracy of any fixed classifier is linear in ${\bm{p}}$, then the core of the proof is a convexity argument: the optimality of $\mathcal{A}^{\mathrm{opt}}$ implies that $\mathrm{acc}_{\bm{p}}(h_{\bm{p}})$ is convex as a pointwise maximum of linear functions, with gradients given by marginal accuracies. From this we identify $f({\bm{p}})=\mathrm{acc}_{\bm{p}}(h_{\bm{p}})$ as a potential function, and leverage the known correspondence between Nash equilibria of population games and local extrema of potential functions. Finally, distinct oracle classifiers imply strict convexity of $f({\bm{p}})$, and the stability of single-population states follows from the fact that convex functions over the simplex are locally-maximized at the vertices, and the accuracy condition follows from the fact that $f({\bm{p}})$ is a Lyapunov function. Full proof appears in Section D.5.
Extensions.
To simplify presentation, the statement of Theorem 1 assumes that the oracle classifiers for each group are distinct. When this assumption doesn’t hold (e.g., when $D_{k}=D_{k^{\prime}}$ for some $k\neq k^{\prime}$), a generalized version of the theorem applies: rather than single-group stable equilibria, dynamics converge towards convex combinations of groups that share an identical oracle classifier with identical marginal accuracy (See Section D.5.1). For heterogeneous fitness where retention functions $\nu_{k}$ may differ across groups, we show that the retention functions we support induce potential games with transformed potential functions. Survival is determined according to the maximizers of the transformed potential function, and the quantity that increases over time becomes the average fitness $\sum_{k}p_{k}\nu_{k}\left(\mathrm{acc}_{k}(h_{\bm{p}})\right)$ (See Section D.5.2). Finally, we also note that a similar characterization also holds for classifiers trained once at the onset (See Section D.4).
@@
Theorem 1 suggests that for repeatedly trained oracle classifiers, natural selection will amplify the dominance of groups having the highest fitness at the onset. From an evolutionary perspective, this reflects the competitive exclusion principle [17], which states that natural selection tends towards exclusive survival of the fittest when multiple species compete over the same resources. Stability implies that once a group dominates, other groups will remain excluded. While this is the general trend, coexistence is nonetheless enabled by retraining, in the form of possible equilibria in which multiple groups survive. That is, there may exist population states ${\bm{p}}^{*}$ such that $p^{*}_{k}>0$ for several groups $k$, and for which this remains to hold when the classifier is retrained. The crux, unfortunately, is that such points are evolutionarily unstable, in the sense that small perturbations around ${\bm{p}}^{*}$ will push dynamics away. This implies that optimal retraining—unconstrained by limited data and compute—induces a strong form of competition between groups. But the existence of mixed equilibria suggests that other algorithms might be able to encourage symbiosis, which we explore next.
5 Avenues for Coexistence
Figure 3: Graphical illustration of theoretical coexistence results. Top row presents data distributions, bottom row shows the corresponding prediction games. (Left) Stable and mutualistic coexistence induced by the use of a proxy loss (Theorem 1). (Center) Stable coexistence induced by interpolation (Section D.7). (Right) Mutualistic coexistence induced by a finite training set (Section D.8).
Intuitively, oracle classifiers drive dynamics away from mixture equilibria because the learning objective is fully aligned with, and so reinforces, self-selection. This motivates the natural question: Can repeated training drive dynamics also towards coexistence, and when? Here we illustrate two ways for symbiosis to arise: as a result of inherent limitations of learning in practice (i.e., limited data or compute), and by applying an ad-hoc, dynamics-aware learning algorithm designed to stabilize outcomes.
A context-dependent view on coexistence.
It is important to emphasize that coexistence in our setting is neither innately ‘good’ or ‘bad’, but rather, its desirability depends on context. Hence, the question of whether coexistence should be encouraged depends on the task at hand, the nature of user choices, and the goals of the system designer. For instance, recommendation platforms may aim to preserve both content and audience diversity, making coexistence desirable. Conversely, in consumer markets having multiple competing standards (e.g., video encoding format, compression schemes, payment platforms), there are benefits to converging on the single best alternative. This context-dependent stance aligns well with the notion of evolution, which in itself is neutral.
5.1 Learning in Practice
While the goal in learning is to minimize the expected 0-1 loss, in practice learning often resorts to optimizing proxy objectives over sampled data via empirical risk minimization. Here we show that, perhaps surprisingly, basic aspects of this approach—namely the use of surrogate losses, access to finite data, and interpolation methods—can create conditions that facilitate mutualism, in which all groups benefit from coexistence. For each aspect, we demonstrate through a carefully crafted example how it can act as a mechanism that enables stable coexistence. Here we focus our discussion on surrogate loss as a representative construction; for finite data and memorization, we illustrate the underlying distribution and induced game in Figure 3, and defer the formal analysis to Section D.6.
Coexistence by surrogate loss.
A common approach to empirical risk minimization is to replace the 0-1 loss with a convex surrogate loss, such as the hinge loss used in the classic Soft-SVM algorithm [9]. This enables optimization, but introduces bias: whereas the 0-1 loss applies uniform penalties, surrogates penalize misclassified points relative to their distance from the decision boundary of $h$. Our next result uses this artifact to construct a game having two groups that are complementary in their incurred bias, and hence balance each other to induce mutualistic coexistence.
Theorem 1.
There exists an evolutionary prediction game in which learning with the hinge loss and $\ell_{2}$ regularization induces a mixed equilibrium that is both stable and fitness-maximizing.
The construction is outlined in Figure 3 (Left), and proof is provided in Section D.6. The proof leverages the bias of hinge loss against minority classes: For each group alone, minimizing the hinge loss biases the learned classifier against the minority class, which is suboptimal. But since each group has a different minority class, these biases cancel out at the mixed equilibrium. The surprising property is that this state is also evolutionarily stable: when one group grows slightly (i.e., $p_{1}=0.5+\epsilon$), it loses more from the fact that the other group shrinks (i.e., $p_{2}=0.5-\epsilon$) than it gains from its own growth. In this sense, each group needs the other to maintain its existence, giving rise to stable and mutualistic symbiosis. Section D.6.3 expands on this idea to show that varying the amount of regularization leads to bifurcations, in which the game transitions between 3, 1, and 5 equilibria.
@@ 5.1 Learning in Practice
In the example game above, as in those we construct for finite data (Section D.8; Figure 3 (Right)) and interpolation (Section D.7; Figure 3 (Center)), mutualism is enabled by how user self-selection compensates for the algorithm’s imperfections. When the resulting bias is complementary across groups, this can give rise to (implicit) cooperation. One interpretation is that the existence of each group acts as regularization on others: from the perspective of group $k$, the objective can be written as $\operatorname*{arg\,max}_{h}\mathrm{acc}_{k}(h)+\lambda R(h)$ where $R(h)=\sum_{j\neq k}\mathrm{acc}_{k}^{\prime}(h)$ is a data-dependent regularizer with coefficient $\lambda=\frac{1}{p_{k}}\sum_{j\neq k}p_{j}$. Nonetheless, our results should not be taken to imply that proxy losses or finite data are a good means to achieve coexistence; rather, they highlight how mutualism can organically materialize when learning affects different groups differently. We explore this further empirically in Section 6.1.